OpenStudy (anonymous):

i just cant wrap my head around it factor x^2-10x+21

OpenStudy (anonymous):

7 and 3!

jimthompson5910 (jim_thompson5910):

Find two numbers that multiply to 21 AND add to -10: These numbers are -7 and -3 x^2-10x+21 x^2-7x-3x+21 x(x-7)-3(x-7) (x-7)(x-3) So x^2-10x+21 = (x-7)(x-3)

OpenStudy (anonymous):

lol my first thoughts were 7 and 3 too XD

OpenStudy (saifoo.khan):

\[\Huge (x-3)(x-7)\]

OpenStudy (anonymous):

ohhhhhhhh!!!!!! two negatives make a positive thanks guys

OpenStudy (anonymous):

what is that cool way that they teach nowadays where the leading and last coefficient = yada yada... i never use that way

OpenStudy (anonymous):

x^2-10x+21 x^2-7x-3x+21 (x-7)(x-3)

OpenStudy (anonymous):

leading and last are only for problems with a coefficent on the first x

OpenStudy (anonymous):

oh hahaha i never use it so i wouldn't know haha

OpenStudy (anonymous):

like 7x^2+5x+30 or something. factors of 7 and factors of 30

OpenStudy (anonymous):

4y^2_49????

OpenStudy (anonymous):

*-

jimthompson5910 (jim_thompson5910):

4y^2-49 = (2y+7)(2y-7) using the difference of squares rule.